OpenStudy (anonymous):

tan^2 x - 3 = 0 for 0 less than or equal to x less than 2pi

OpenStudy (anonymous):

whats ur question?

OpenStudy (anonymous):

my question is tan^2 x - 3 = 0 for 0 less than or equal to x less than 2pi

OpenStudy (anonymous):

right. write tan^2(x) interms of cos2x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok im not getting this

OpenStudy (anonymous):

first write tan^2(x) interms of cos(2x)

OpenStudy (anonymous):

ok im not getting this

OpenStudy (anonymous):

ok im not getting this

OpenStudy (anonymous):

ok im not getting this

OpenStudy (anonymous):

\[\tan^2(x) = \frac{ 1-\cos(2x) }{ 1+\cos(2x)}\] use this formula and find the value of cos(2x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

umm can you explain it more?

OpenStudy (anonymous):

\[\tan^2(x)=\frac{1−\cos(2x)}{1+\cos(2x)}\] so \[\frac{1−\cos(2x)}{1+\cos(2x)} = 3\] now find cos(2x)

OpenStudy (anonymous):

im still not getting this :/

OpenStudy (anonymous):

@CliffSedge do you know how i can do this problem?

OpenStudy (anonymous):

cross multiply and simplify. find cos(2x) you should get cos(2x) = -1/2

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

I'll take a look at it.

OpenStudy (anonymous):

Everything chenna showed looks fine. What part are you still stuck on @onegirl ?

OpenStudy (anonymous):

nvm i think i got it

OpenStudy (anonymous):

x = 0 to 2pi right?

OpenStudy (anonymous):

thats the answer right?

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

no. that is not the answer. you need to find the value of x. did you get cos(2x) = -1/2 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so x = 3pi/2, pi/6, and 5pi/6?

OpenStudy (anonymous):

no. when does cos(t) = -1/2 ?

OpenStudy (anonymous):

when i crossed multiply

OpenStudy (anonymous):

I meant for which value of t, cos(t) = -1/2

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

i think i got it thanks.

OpenStudy (anonymous):

so what is the value of x you got?

OpenStudy (anonymous):

x = pi/4 ,3pi/4, ,5pi/4, 7pi/4

OpenStudy (anonymous):

no. those values are wrong. when does the value of cosine function equals to 1/2 ??

OpenStudy (anonymous):

ok